# On the non-Markovian quantum control dynamics

### v3

@@ Abstract
In this paper, we study both open-loop control and closed-loop measurement feedback control of non-Markovian quantum dynamics resulting from the interaction between a quantum system and its environment. We use the widely studied cavity quantum electrodynamics (cavity-QED) system as an example, where an atom interacts with the environment composed of a collection of oscillators. In this scenario, the stochastic interactions between the atom and the environment can introduce non-Markovian characteristics into the evolution of quantum states, differing from the conventional Markovian dynamics observed in open quantum systems. As a result, the atom’s decay rate to the environment varies with time and can be described by nonlinear equations. The solutions to these nonlinear equations can be analyzed in terms of the stability of a nonlinear control system. Consequently, the evolution of quantum state amplitudes follows linear time-varying equations as a result of the non-Markovian quantum transient process. Additionally, by using measurement feedback through homodyne detection of the cavity output, we can modulate the steady atomic and photonic states in the non-Markovian process. When multiple coupled cavity-QED systems are involved, measurement-based feedback control can influence the dynamics of high-dimensional quantum states, as well as the resulting stable and unstable subspaces.
Index Terms:
quantum non-Markovian dynamics, open quantum system, quantum open-loop control, quantum measurement feedback control.
@@ I Introduction
Quantum control has garnered much attention due to its potential applications in quantum optics [1, 2], quantum information processing [3], quantum engineering [4] and others [5]. Control methods for quantum systems can be categorized into the open-loop control and closed-loop control, similar to that in classical systems [1]. In open-loop control, designed or iteratively optimized control pulses are utilized to produce required gate operations in quantum computations [6, 7]. When a control field is applied upon an atom, the atom can be excited, leading to the generation of single or multiple photons for quantum networking [8]. On the other hand, closed-loop quantum control can be realized through coherent feedback or measurement feedback methods. For instance, when an atom or cavity quantum electrodynamics (cavity-QED) system is coupled with a waveguide, a coherent feedback channel can be established using photons transmitted in the waveguide [9]. Subsequently, the atomic dynamics and photonic states can be modulated by tuning the parameters of the coherent feedback loop [10, 9, 11]. This form of coherent feedback dynamics can be represented as a linear control system with delays determined by the length of the feedback loop [10, 9]. Then the quantum state dynamics can be interpreted in terms of the stability of a linear control system with delays [12, 11, 13].
@@
In the realm of open quantum system control, it is typically assumed that the environmental evolution timescale is considerably shorter than that of the quantum system of interest, and the decoherence rate of the quantum control system to the environment can be simplified as static [21]. This assumption allows for the modeling of the interaction between the quantum system and the environment using a master equation with a static decaying rate, which is a widely employed Markovian approximation method for studying quantum dynamics in open systems. However, in numerous scenarios, this Markovian approximation proves inadequate for comprehensively analyzing the dynamics of open quantum systems [21, 22]. For instance, in the experimental implementation utilizing nuclear magnetic resonance (NMR) [23], where the NMR qubit interacts with a non-Markovian environment characterized by a randomized configuration of modulated radio-frequency fields. In such setups, the interaction between the NMR and the environment can lead to information backflow from the environment to the qubit, thus the decoherence of the qubit can be nonmonotonic, which is different from the monotonic decoherence of a qubit in the Markovian environment [24]. Moreover, experimental evidence of information backflow induced by non-Markovianity has also been observed in optical systems [25]. Given that traditional quantum control strategies relying on Markovian approximation are not directly applicable in these instances, it is required to explore the open-loop and closed-loop quantum control dynamics for non-Markovian open quantum systems. Additionally, it is worth noting that the traditional Morkovian scenario can be encompassed within the broader framework of non-Markovian settings as a simplified special case.
@@
The interaction between a quantum system and its environment, characterized as non-Markovian, can be effectively represented by the integral stochastic Schrödinger equation [26]. This approach incorporates the influence of the environment on quantum states through a complex-valued stochastic process with an integral kennel relative with historic memory effects in the time domain [27, 26]. Alternatively, after averaging the quantum states over the environmental noise, the non-Markovian dynamics can be described using a master equation featuring time-varying Lindblad components [28, 29], resulting from the memory effect inherent in non-Markovian dynamics [30]. Consequently, various non-Markovian quantum control techniques have been developed including the learning [31, 32] or hybrid control approaches upon quantum systems and the modeled classical non-Markovian environment [33], expanding upon the foundation laid by Markovian quantum control methods [34, 35, 36].
@@
In this paper, we utilize the commonly employed cavity-QED system depicted in Fig. 1 as an example to investigate the dynamics of non-Marovkovian quantum control in a novel manner. We for the first time clarify that, the time-varying parameters indicating a non-Markovian process for open quantum system can be analyzed from the perspective of nonlinear process. Based on this, we illustrate the transition from non-Markovian transient dynamics to steady Markovian dynamics using stability theory in nonlinear control. Subsequently, we describe the evolution of atomic and photonic states in the non-Markovian quantum dynamics using linear time varying (LTV) control equations, where external drives or the measurement feedback can influence the quantum dynamics. Lastly, we extend our analysis to encompass non-Markovian dynamics in systems where atoms in multiple coupled cavities interact with a non-Markovian environment. Then the quantum states can be categorized into the stable and unstable subspaces based on the application of measurement feedback controls.
@@ II Non-Markovian quantum control for open systems
In this section, we introduce the quantum control modeling based on the non-Markovian interactions between a cavity-QED system and its environment. The model is made up of the linear time varying dynamics for the interactions between a multi-level atom and a cavity, and the non-Markovian decaying of the quantum system to the environment governed by nonlinear dynamics.
@@
$H=\omega_{c}a^{{\dagger}}a+\sum_{\omega}\omega b_{\omega}^{{\dagger}}b_{\omega}+\sum_{n=0}^{N-1}\omega_{n}|n\rangle\langle n|+H_{t},$ (1)
@@
where $\omega_{c}$ is the resonant frequency of the cavity with the annihilation (creation) operator $a$ ($a^{{\dagger}}$), the second component represents a group of oscillators in the environment modeled by the annihilation (creation) operator $b_{\omega}$ ($b_{\omega}^{{\dagger}}$) with different frequencies $\omega$, the third component is the atom Hamiltonian with $\omega_{n}$ represents the energy of the $n$-th level represented with the multiplication of the vector $|n\rangle$ and $\langle n|$, $H_{t}$ is composed with the interaction Hamiltonian between the atom and cavity, as well as the interaction Hamiltonian between the atom and environment.
@@
For the $N$-level atom coupled with the cavity, as in Fig. 1, the Hamiltonian $H_{t}$ in Eq. (1) can be equivalently represented in the interaction picture as [9, 37]
$\bar{H}_{t}=$ $\sum_{n=1}^{N-1}g_{n}\left(e^{-i\Delta_{n}t}\sigma^{-}_{n}a^{{\dagger}}+e^{i\Delta_{n}t}\sigma^{+}_{n}a\right)$ (2)
$+\sum_{n=1}^{N-1}\sum_{\omega}\chi_{\omega}^{(n)}\left(\sigma^{-}_{n}b_{\omega}^{{\dagger}}+\sigma^{+}_{n}b_{\omega}\right),$
@@
where we denote $\tilde{\omega}_{n}=\omega_{n}-\omega_{n-1}$, $\Delta_{n}=\tilde{\omega}_{n}-\omega_{c}$ is the detuning between the cavity resonant frequency and the transition frequency of two neighborhood atom energy levels, $g_{n}$ represents the coupling strengths between the cavity and different atom energy levels, the lowering operator $\sigma^{-}_{n}=|n-1\rangle\langle n|$, and the rising operator $\sigma^{+}_{n}=|n\rangle\langle n-1|$. The first part on RHS represents the interaction between the atom and cavity, $\sigma^{-}_{n}a^{{\dagger}}$ represents that when the atom decays from the $n$-th energy level to the $(n-1)$-th energy level, it can emit one photon into the cavity. The following Hermitian conjugate $\sigma^{+}_{n}a$ represents that the atom can absorb one photon from the cavity and be excited to a higher energy level. The second part on RHS similarly represents the interaction between the multi-level atom and environment, and the coupling strength between the $n$-th energy level and the environmental oscillator with the frequency $\omega$ is $\chi_{\omega}^{(n)}$. $\sigma^{-}_{n}b_{\omega}^{{\dagger}}$ represents that an emitted filed generated by the atom’s decay from the $n$-th to $(n-1)$-th energy level can be absorbed by the environment, similar for the following Hermitian conjugate. For convenience, we denote $\tilde{g}_{n}=g_{n}e^{i\Delta_{n}t}$ in the following.
@@
The different energy levels of the ladder type atom can be coupled to the environment with the operator [38, 39, 40]
$L=\sum_{n=1}^{N-1}\sqrt{\kappa_{n}}L_{n}=\sum_{n=1}^{N-1}\sqrt{\kappa_{n}}|n-1\rangle\langle n|,$ (3)
@@
where $\kappa_{n}$ represents the decaying rate of the atomic state $|n\rangle$ to the environment and $L_{n}=|n-1\rangle\langle n|$. More details are introduced in Appendix A.
@@
The dynamics of the above quantum system can be modeled with the Schrödinger equation $\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=-iH|\psi(t)\rangle$ , and the components for the interaction between the quantum system and the environment can be written as a stochastic format, as introduced in Refs. [27, 26, 41, 42]. Combined with Eq. (3), the Schrödinger equation can be equivalently written as  [27, 26, 41, 42]
$\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=$ $-i\bar{H}|\psi(t)\rangle$ (4)
@@
$+\left[L|\psi(t)\rangle z_{t}-L^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi(t)\rangle}{\delta z_{s}}\mathrm{d}s\right],$
@@
where $\bar{H}$ represents the interaction component between the atom and cavity in Eq. (2), $z_{t}$ is a complex stochastic process determined by the environment with $E(z_{t}^{*}z_{s})=\alpha(t,s)$, $E(z_{t}z_{s})=0$, and
@@
where $\gamma^{-1}$ represents the environmental memory time scale and $\Omega$ represents the environmental central frequency for the modeled oscillators [27].
Generalized from the definition of a quantum Markov process in [43], the quantum dynamics governed by equation (4) is non-Markovian when $\gamma$ is finite, as it includes a memory term wherein the quantum state $|\psi(t)\rangle$ depends on past noise $z_{s}$, weighted by a nonzero kernel $\alpha(t,s)$ for $s<t$ [26].
Alternatively, based on Eq. (4), the quantum dynamics can be equivalently represented by the density matrix $\rho(t)=|\psi(t)\rangle\langle\psi(t)|$, whose dynamics is governed by the following master equation according to the derivation in Appendix B,
$\dot{\rho}=$ $-i[\bar{H},\rho]+\sum_{n=1}^{N-1}\left[F_{n}(t)+F_{n}^{*}(t)\right]L_{n}\rho L_{n}^{{\dagger}}$ (6)
$-\sum_{n=1}^{N-1}\left[F_{n}(t)L_{n}^{{\dagger}}L_{n}\rho+F_{n}^{*}(t)\rho L_{n}^{{\dagger}}L_{n}\right],$
@@
and the integration in Eq. (7) represents the memory effect of the non-Markovian process based on the interaction between the atom and environment.
The reduced quantum Markovian dynamics can be distinguished according to the following remark based on the integral kernel $\alpha(t,s)$.
@@
When $\gamma\rightarrow\infty$, the interaction between the system and the environment is Markovian and $\alpha(t,s)=\delta(t-s)$ [27, 43]. Here,
$F_{n}(t)=\int_{0}^{t}\delta(t-s)f_{n}(t,s)\mathrm{d}s=\frac{1}{2}f_{n}(t,t)\triangleq\frac{\chi_{n}}{2},$ (8)
@@
II-A Single multi-level atom coupled with the cavity
According to the non-Markovian master equation (6), the dynamics of the mean-value of an arbitrary operator O is governed by $\dot{\langle\textbf{O}\rangle}={\rm Tr}\left[\dot{\rho}\textbf{O}\right]$ [15]. To clarify the atomic dynamics, we study the dynamics of the operator $\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$ representing the population or probability that the atom is excited at the $n$-th energy level. Because of the coupling between the atom and the cavity, the dynamics of $\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$ is affected by the exchange of energy between the atom and the cavity, which can be evaluated by the dynamics of the operator $\langle\sigma^{+}_{n}a\rangle$, representing that the atom can absorb one photon from the cavity to be excited to a higher energy level, or the reverse process evaluated by $\langle\sigma^{-}_{n}a^{{\dagger}}\rangle$. Besides, the number of photons in the cavity can be evaluated with the operator $\langle a^{{\dagger}}a\rangle$ [44]. Then we can derive the equation of the mean values of the operators with $n=1,2,\dots,N-1$ as [45, 46, 47, 48]
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=-i\tilde{g}_{n}\langle\sigma^{+}_{n}a\rangle+i\tilde{g}_{n}^{*}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle+i\tilde{g}_{n+1}\langle\sigma^{+}_{n+1}a\rangle$
$~{}~{}~{}~{}~{}~{}-i\tilde{g}_{n+1}^{*}\langle\sigma^{-}_{n+1}a^{{\dagger}}\rangle-\left(F_{n}(t)+F_{n}^{*}(t)\right)\kappa_{n}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}+\left(F_{n+1}(t)+F_{n+1}^{*}(t)\right)\kappa_{n+1}\langle\sigma^{+}_{n+1}\sigma^{-}_{n+1}\rangle,$ (9a)
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}a\rangle=-i\Delta_{n}\langle\sigma^{+}_{n}a\rangle-i\tilde{g}_{n}^{*}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+i\tilde{g}_{n}^{*}\langle a^{{\dagger}}a\rangle-F_{n}^{*}(t)\kappa_{n}\langle\sigma^{+}_{n}a\rangle,$ (9b)
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle=i\Delta_{n}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle+i\tilde{g}_{n}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-i\tilde{g}_{n}\langle a^{{\dagger}}a\rangle-F_{n}(t)\kappa_{n}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle,$ (9c)

### v4

@@ Abstract
In this paper, we study both open-loop control and closed-loop measurement feedback control of non-Markovian quantum dynamics arising from the interaction between a quantum system and its environment. We use the widely studied cavity quantum electrodynamics (cavity-QED) system as an example, where an atom interacts with the environment composed of a collection of oscillators. In this scenario, the stochastic interactions between the atom and the environment can introduce non-Markovian characteristics into the evolution of quantum states, differing from the conventional Markovian dynamics observed in open quantum systems. As a result, the atom’s decay rate to the environment varies with time and can be described by nonlinear equations. The solutions to these nonlinear equations can be analyzed in terms of the stability of a nonlinear system. Consequently, the evolution of quantum state amplitudes follows linear time-varying equations as a result of the non-Markovian quantum transient process. Additionally, by using measurement feedback through homodyne detection of the cavity output, we can modulate the steady atomic and photonic states in the non-Markovian process. When multiple coupled cavity-QED systems are involved, measurement-based feedback control can influence the dynamics of high-dimensional quantum states, as well as the resulting stable and unstable subspaces.
@@ I Introduction
Quantum control has garnered much attention due to its potential applications in quantum optics [1, 2], quantum information processing [3], quantum engineering [4] and others [5]. Control methods for quantum systems can be categorized into the open-loop control and closed-loop control, similar to that in classical systems [1]. In open-loop control, designed or iteratively optimized control pulses are utilized to produce required gate operations in quantum computations [6, 7]. When a control field is applied upon an atom, the atom can be excited, leading to the generation of single or multiple photons for quantum networking [8]. On the other hand, closed-loop quantum control can be realized through coherent feedback or measurement feedback methods. For instance, when an atom or cavity quantum electrodynamics (cavity-QED) system is coupled to a waveguide, a coherent feedback channel can be established using photons transmitted in the waveguide [9]. Subsequently, the atomic dynamics and photonic states can be modulated by tuning the parameters of the coherent feedback loop [10, 9, 11]. This form of coherent feedback dynamics can be represented as a linear control system with delays determined by the length of the feedback loop [10, 9]. Then the quantum state dynamics can be interpreted in terms of the stability of a linear control system with time delays [11, 12].
@@
In the realm of open quantum system control, it is typically assumed that the environmental evolution timescale is considerably shorter than that of the quantum system of interest, and the decoherence rate of the quantum control system to the environment can be simplified as static [20]. This assumption allows for the modeling of the interaction between the quantum system and the environment using a master equation with a static decay rate, which is a widely employed Markovian approximation method for studying quantum dynamics in open systems. However, in numerous scenarios, this Markovian approximation proves inadequate for comprehensively analyzing the dynamics of open quantum systems [20, 21]. For instance, in the experimental implementation utilizing nuclear magnetic resonance (NMR) [22], the NMR qubit interacts with a non-Markovian environment characterized by a randomized configuration of modulated radio-frequency fields. In such setups, the interaction between the NMR and the environment can lead to information backflow from the environment to the qubit, thus the decoherence of the qubit can be nonmonotonic, which is different from the monotonic decoherence of a qubit in the Markovian environment [23]. Moreover, experimental evidence of information backflow induced by non-Markovianity has also been observed in optical systems [24]. Due to the existence of non-Markovian noises in quantum information processing, quantum error correcting protocols based on three- or five-qubit codes [25], and Petz map [26] can be used to enhance the performance against non-Markovian noises. Relied on these protocols, and generalized from the Markovian scenario [18], measurement feedback control based on non-Markovian quantum dynamical systems can be potentially applied in quantum error corrections [27]. Given that traditional quantum control strategies within Markovian approximations are not directly applicable in these instances, it is required to explore the open-loop and closed-loop quantum control dynamics for non-Markovian open quantum systems. Additionally, it is worth noting that the traditional Morkovian scenario can be encompassed within the broader framework of non-Markovian settings as a simplified special case.
@@
The interaction between a quantum system and its environment, characterized as non-Markovian, can be effectively represented by the integral stochastic Schrödinger equation [28]. This approach incorporates the influence of the environment on quantum states through a complex-valued stochastic process with an integral kennel relative with historic memory effects in the time domain [29, 28]. Alternatively, after averaging the quantum states over the environmental noise, the non-Markovian dynamics can be described using a master equation featuring time-varying Lindblad components [30, 31], resulting from the memory effect inherent in non-Markovian dynamics [32]. Consequently, various non-Markovian quantum control techniques have been developed including the learning [33, 34] and hybrid control approaches upon the quantum system and environment [35], which are generalized from Markovian quantum control methods [36, 37, 38].
@@
In this paper, we utilize the commonly employed cavity-QED system depicted in Fig. 1 as an example to investigate the dynamics of non-Marovkovian quantum control in a novel manner. We clarify that, the time-varying parameters indicating a non-Markovian process for open quantum system can be analyzed from the perspective of nonlinear process. Based on this, we illustrate the transition from non-Markovian transient dynamics to steady Markovian dynamics using the stability theory in nonlinear control. Subsequently, we describe the evolution of atomic and photonic states in the non-Markovian quantum dynamics with linear time varying (LTV) control equations, where external drives or the measurement feedback can influence the quantum dynamics. Lastly, we extend our analysis to encompass non-Markovian dynamics in systems where atoms in multiple coupled cavities interact with a non-Markovian environment. Then the quantum states can be categorized into the stable and unstable subspaces based on the application of measurement feedback controls.
@@ II Non-Markovian quantum control for open systems
In this section, we introduce the quantum control modeling based on the non-Markovian interactions between a cavity-QED system and its environment. The model is made up of the linear time varying dynamics for the interactions between a multi-level atom and a cavity, and the non-Markovian decay of the quantum system to the environment governed by nonlinear dynamics.
@@
$H=\omega_{c}a^{{\dagger}}a+\sum_{\omega}\omega b_{\omega}^{{\dagger}}b_{\omega}+\sum_{n=0}^{N-1}\omega_{n}|n\rangle\langle n|+H_{I},$ (1)
@@
where $\omega_{c}$ is the resonant frequency of the cavity with the annihilation (creation) operator $a$ ($a^{{\dagger}}$), the second component represents a group of oscillators in the environment modeled by the annihilation (creation) operator $b_{\omega}$ ($b_{\omega}^{{\dagger}}$) with different frequencies $\omega$, the third component is the atom Hamiltonian with $\omega_{n}$ represents the energy of the $n$-th level represented with the multiplication of the vector $|n\rangle$ and $\langle n|$, $H_{I}$ is composed with the interaction Hamiltonian between the atom and cavity, as well as the interaction Hamiltonian between the atom and environment.
@@
For the $N$-level atom coupled to the cavity, as in Fig. 1, the Hamiltonian $H_{I}$ in Eq. (1) can be equivalently represented in the interaction picture as [9, 39]
$\bar{H}_{I}=$ $\sum_{n=1}^{N-1}g_{n}\left(e^{-i\Delta_{n}t}\sigma^{-}_{n}a^{{\dagger}}+e^{i\Delta_{n}t}\sigma^{+}_{n}a\right)+\sum_{n=1}^{N-1}\sum_{\omega}\chi_{\omega}^{(n)}\left(\sigma^{-}_{n}b_{\omega}^{{\dagger}}+\sigma^{+}_{n}b_{\omega}\right),$ (2)
@@
where we denote $\tilde{\omega}_{n}=\omega_{n}-\omega_{n-1}$, $\Delta_{n}=\tilde{\omega}_{n}-\omega_{c}$ is the detuning between the cavity resonant frequency and the transition frequency of two neighborhood atom energy levels, $g_{n}$ represents the coupling strengths between the cavity and different atom energy levels, the lowering operator $\sigma^{-}_{n}=|n-1\rangle\langle n|$, and the rising operator $\sigma^{+}_{n}=|n\rangle\langle n-1|$. The first part on the right-hand side (RHS) represents the interaction between the atom and cavity, $\sigma^{-}_{n}a^{{\dagger}}$ represents that when the atom decays from the $n$-th energy level to the (n1)(n-1)-th energy level, it can emit one photon into the cavity. The following Hermitian conjugate $\sigma^{+}_{n}a$ represents that the atom can absorb one photon from the cavity and be excited to a higher energy level. The second part on RHS similarly represents the interaction between the multi-level atom and environment, and the coupling strength between the $n$-th energy level and the environmental oscillator with the frequency $\omega$ is $\chi_{\omega}^{(n)}$. $\sigma^{-}_{n}b_{\omega}^{{\dagger}}$ represents that an emitted filed generated by the atom’s decay from the $n$-th to (n1)(n-1)-th energy level can be absorbed by the environment, similar for the following Hermitian conjugate. For convenience, we denote $\tilde{g}_{n}=g_{n}e^{i\Delta_{n}t}$ in the following.
@@
Different energy levels of the ladder type atom can be coupled to the environment with the operator [40, 41, 42]
$L=\sum_{n=1}^{N-1}L_{n}=\sum_{n=1}^{N-1}\sqrt{\kappa_{n}}|n-1\rangle\langle n|,$ (3)
@@
where $\kappa_{n}$ represents the decay rate of the atomic state $|n\rangle$ to the environment and $L_{n}=\sqrt{\kappa_{n}}|n-1\rangle\langle n|$. More details are introduced in Appendix A.
@@
The dynamics of the above quantum system can be modeled with the Schrödinger equation $|\dot{\psi}(t)\rangle=-iH|\psi(t)\rangle$, and the influence by the environmental oscillators can be modeled as a stochastic process with Markovian or non-Markovian statistical properties, as introduced in Refs. [29, 28, 43, 44]. Due to Eq. (3), the Schrödinger equation can be written as  [29, 28, 43, 44]
@@
$\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=$ $-i\bar{H}|\psi(t)\rangle+\left[L|\psi(t)\rangle z_{t}-L^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi(t)\rangle}{\delta z_{s}}\mathrm{d}s\right],$ (4)
@@
where $\bar{H}$ represents the interaction component between the atom and cavity in Eq. (2), $z_{t}$ is a complex stochastic process determined by the environment with $E(z_{t}^{*}z_{s})=\alpha(t,s)$, $E(z_{t}z_{s})=0$, $\delta|\psi(t)/\delta z_{s}$ is a functional derivative of $|\psi(t)\rangle$ with respect to $z_{s}$, and
@@
where $\gamma^{-1}$ represents the environmental memory time scale and $\Omega$ represents the environmental central frequency for the modeled oscillators [29]. Generalized from the definition of a quantum Markov process in [45], the quantum dynamics governed by equation (4) is non-Markovian when $\gamma$ is finite, as it includes a memory term wherein the quantum state $|\psi(t)\rangle$ depends on past noise $z_{s}$, weighted by a nonzero kernel $\alpha(t,s)$ for $s<t$ [28].
Alternatively, the quantum dynamics in Eq. (4) can be equivalently represented by the density matrix $\rho(t)=|\psi(t)\rangle\langle\psi(t)|$, whose dynamics is governed by the following master equation according to the derivation in Appendix B,
$\dot{\rho}=$ $-i[\bar{H},\rho]+\sum_{n=1}^{N-1}\left[F_{n}(t)+F_{n}^{*}(t)\right]L_{n}\rho L_{n}^{{\dagger}}-\sum_{n=1}^{N-1}\left[F_{n}(t)L_{n}^{{\dagger}}L_{n}\rho+F_{n}^{*}(t)\rho L_{n}^{{\dagger}}L_{n}\right],$ (6)
@@
$f_{n}(t,s)$ is a function to be determined according to the memory kernel $\alpha(t,s)$ [29]. We denote the reduced value $\chi_{n}=f_{n}(t,t)$ when $t=s$, and $\chi_{n}$ can be different constants for different channels indexed by $n$ [29]. Then due to Eq. (77) in Appendix A, the trajectory of $F_{n}(t)$ can equivalently rely on $\chi_{n}$ rather than $f_{n}(t,s)$. For the $n$-th Lindbladian in Eq. (6), we briefly introduce the physical meanings in Table I.
TABLE I: Physical meanings of the components in Eq. (6)
Index Decaying operator Integral kernel Interaction rate
$n$ $L_{n}$ $\alpha(t,s)$ $F_{n}(t)$
The reduced Markovian case is distinguished as follows.
@@
When $\gamma\rightarrow\infty$, the interaction between the quantum system and environment is Markovian with $\alpha(t,s)=\delta(t-s)$, then [29, 45]
$F_{n}(t)=\int_{0}^{t}\delta(t-s)f_{n}(t,s)\mathrm{d}s=\frac{\chi_{n}}{2},$ (8)
@@
II-A Single multi-level atom coupled to the cavity
Based on the non-Markovian Schrödinger equation (4) in a vector format, or the non-Markovian master equation (6) in a matrix format, the non-Markovian atomic dynamics can be evaluated by the mean-value of an arbitrary operator O is governed by $\dot{\langle\textbf{O}\rangle}={\rm Tr}\left[\dot{\rho}\textbf{O}\right]$ [14]. For example, the operator mean value $\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$ represents the population or probability that the atom is excited at the $n$-th energy level. Because of the coupling between the atom and the cavity, the dynamics of $\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$ is affected by the exchange of energy between the atom and the cavity, which can be evaluated by the dynamics of the operator mean value $\langle\sigma^{+}_{n}a\rangle$, representing that the atom can absorb one photon from the cavity to be excited to a higher energy level, or the reverse process evaluated by $\langle\sigma^{-}_{n}a^{{\dagger}}\rangle$. Besides, the number of photons in the cavity can be evaluated by the mean value $\langle a^{{\dagger}}a\rangle$ [46]. Then we can derive the equation of the mean values of the operators as [47, 48, 49, 50]
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=-i\tilde{g}_{n}\langle\sigma^{+}_{n}a\rangle+i\tilde{g}_{n}^{*}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle+i\tilde{g}_{n+1}\langle\sigma^{+}_{n+1}a\rangle-i\tilde{g}_{n+1}^{*}\langle\sigma^{-}_{n+1}a^{{\dagger}}\rangle$
$~~~~~~~~~~~~~~~~-\left(F_{n}(t)+F_{n}^{*}(t)\right)\kappa_{n}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle+\left(F_{n+1}(t)+F_{n+1}^{*}(t)\right)\kappa_{n+1}\langle\sigma^{+}_{n+1}\sigma^{-}_{n+1}\rangle,$ (9a)
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}a\rangle=-i\Delta_{n}\langle\sigma^{+}_{n}a\rangle-i\tilde{g}_{n}^{*}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle+i\tilde{g}_{n}^{*}\langle a^{{\dagger}}a\rangle-F_{n}^{*}(t)\kappa_{n}\langle\sigma^{+}_{n}a\rangle,$ (9b)
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle=i\Delta_{n}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle+i\tilde{g}_{n}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle-i\tilde{g}_{n}\langle a^{{\dagger}}a\rangle-F_{n}(t)\kappa_{n}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle,$ (9c)